# Instructions for *ACL Proceedings

### v1

@@ Appendix B Implementation Details
B.1 Judgment Thinking Initialization
@@
We leverage the prompt templates in Figure 11 and 13 for constructing thinking trace based on proprietary thinking models. The temperature is set as 0.0 to enable more accurate judgment. We also present an example of thinking trace calibration in 10, which can ensure training efficiency and stability.
@@
B.2 Judgment Thinking Optimization
For judgment thinking optimization, we leverage the prompts in Figure 11 and 12 for offline learning, and the prompts in Figure 13 and 14 for online learning. The offline learning is implemented based on Llama-Factory Zheng et al. (2024), and online learning is implemented based on verl Sheng et al. (2024). We present the hyper-parameter settings in Table 10 and 11.

### v2

@@ Appendix B Implementation Details
B.1 Thinking Trace Clipping
As we have explained in Section 3.1, despite the superior reasoning capability of Deepseek-R1, the thinking trace generated are excessively long and could possibly cause difficulty for further optimization and increase computational overhead. Specifically, for our case, we hypothesize that general LLM judgment does not necessarily require a detailed, step-by-step reasoning process. On the contrary, identifying one key factor is sufficient to make correct judgment in most cases. Overly long reasoning traces can cause the model to consider factors that do not affect the preference outcomes, thereby degrading the effectiveness.
Therefore, we introduce Trace Clipping to enhance the effectiveness of thinking initialization, as shown in the Example of 9. Specifically, we observed that the CoTs generated by R1 are composed of two parts: a lengthy reasoning process, and an explanation which summarizes the reasoning process. Therefore, we remove the first part and use the second part as the output trace. Models trained with the clipped trace will also generate the reasoning trace in a way that directly addresses the key points. This not only facilitates subsequent optimization but also enhances the judgment readability.
The experiment analysis of the effectiveness of thinking trace clipping is presented in Appendix A.2.
B.2 Judgment Thinking Initialization
@@
We leverage the prompt templates in Figure 11 and 13 for constructing thinking trace based on proprietary thinking models. The temperature is set as 0.0 to enable more accurate judgment.
@@
B.3 Judgment Thinking Optimization
For judgment thinking optimization, we leverage the prompts in Figure 11 and 12 for offline learning, and the prompts in Figure 13 and 14 for online learning. The offline learning is implemented based on Llama-Factory (Zheng et al. 2024), and online learning is implemented based on verl (Sheng et al. 2024). We present the hyper-parameter settings in Table 17 and 18.
B.4 Criteria for Determining Preference Strength
As we have explained in Section 3.2, for Rule-based Online Learning, we propose an $r_{strength}$ for assessing preference strength, which is defined as the degree to which the judge favors one response over another. In this section, we would like to detail how we get the strength annotation for HelpSteer2-Pref and HH-RLHF datasets.
For the HelpSteer2 dataset, we directly adopt the original data annotations, which were derived from professional human annotators and provide three distinct strength levels.
For the HH-RLHF dataset, we derive strength annotations following the method of (Wang et al. 2024b), which involves:
  1. 1.
    Training multiple Bradley-Terry classifiers on preference data.
  2. 2.
    Averaging the scores assigned to each data point by these classifiers.
  3. 3.
    Categorizing the strength into three tiers (1, 2, 3) based on the score differences between preference pairs.
We opted for a three-level categorization of strength—representing ”slightly better,” ”better,” and ”much better”—due to its clear interpretability. We found that further granularity did not improve results and, in some cases, led to the model ”hacking” to extreme values, as detailed in Appendix B.1.